# Calibrated Physics-Informed Uncertainty Quantification

### v1

@@ 4.3 Differential Operator: Finite-Difference Stencils as Convolutional Kernels
Calibrating neural PDEs using PRE nonconformity scores requires frequent evaluations of the composite differential operator $D$ in Equation 1. For PDEs, this involves estimating numerous spatio-temporal gradients across the discretised domain, ranging from millions in simple cases to billions of gradient operations for complex physics. To address this computational challenge, we developed a scalable gradient estimation method for evaluating physics residual error.
We employ convolution operations with Finite Difference (FD) stencils as convolutional kernels for gradient estimation (Actor et al., 2020; Chen et al., 2024a, b). For instance, the 2D Laplacian operator $\nabla^{2}$, using a central difference scheme with discretisation $h$, can be approximated by
@@
and used as a kernel. This approach is justified by the mathematical equivalence of FD approximations and discrete convolutions. Both represent matrix-vector multiplications of a block Toeplitz matrix with a field vector (Strang, 1986; Fiorentino & Serra, 1991). The efficiency of this method stems from the optimised implementation of convolution operations in machine learning libraries like PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2015).  [^1] The Basic linear Algebra subroutines (BLAS) within these libraries leverage vectorisation and efficient memory access, resulting in significant performance improvements. Our experiments show a 1000x speed-up using torch.nn.functional.conv3d compared to a numpy implementation of the equivalent FD approximation on a standard CPU.
The FD approximation offers several advantages over Automatic Differentiation (AD) for our application. It is compatible with CP as a post-hoc measure, requires no architectural modifications, and is model-agnostic. Furthermore, FD implemented via convolutions is more memory-efficient than AD, which requires storing the entire computational graph. Our focus on the (mis)alignment of neural PDEs with Equation 1 allows us to disregard boundary conditions in our error bar estimations, though we do a preliminary exploration in Appendix E. Utilising finite difference schemes, the residuals are estimated up to the truncation errors of the Taylor approximation (MacKinnon & Johnson, 1991). Though the discretisation of the problem plays a role in the width of the error bars, PRE-CP still guarantees coverage, and this is further explored in Section D.1.
@@ 5 Experiments
Figure 3: Validation plots demonstrating coverage guarantee detailed in Equation 5 obtained by performing CP using PRE across experiments. The average empirical coverage obtained by CP is given on the $y$-axis (ranging from 0 to 1, with 1 representing 100% coverage), while the coverage for which we calibrate is represented on the $x$-axis. We obtain guaranteed coverage while using marginal-CP formulation and near-to-ideal coverage for the joint-CP formulation.
PRE-CP experiments comprise two campaigns. First, we benchmark PRE-CP within standard neural PDEs (Section 5.1 to 5.3). The calibration process (Figure 2) involves: (a) sampling model inputs, (b) calculating PRE(s) scores, and (c) calibrating physical error using marginal and joint-CP formulations. Validation uses the same PDE condition bounds as calibration. This campaign demonstrates our method’s superior computational efficiency and guaranteed coverage versus other Neural-PDE UQ measures (Appendix C).
The second campaign (Section 5.4, 5.5) applies PRE-CP to fusion applications. We enhance tokamak plasma behaviour surrogate models to identify erroneous dispersion regions (Section 5.4) and integrate PRE-CP with tokamak design surrogates to identify viable designs and areas needing additional simulations (Section 5.5). This campaign demonstrates the utility of PRE-CP in complex, practical applications. Reproducible One-dimensional PDE experiments demonstrating PRE-CP are demonstrated in Appendix F.
@@ 5.1 Wave Equation
Figure 4: Wave: (From left to right) neural PDE (FNO) prediction at the last time instance, physics residual error of the prediction, Upper error bars obtained by performing marginal-CP and joint-CP respectively (90% coverage). For brevity, we have only shown the upper error bars of the symmetric prediction sets. $mod$ represents the modulation function in Table 1. The physical inconsistencies within the residual space of the prediction are calibrated and bounded using PRE-CP.
@@
We solve Equation 10 within the domain $x,y\in[-1,1]$, $t\in[0,1.0]$ with $c=1.0$ using a spectral solver with periodic boundary conditions (Canuto et al., 2007). The initial conditions are parameterised by the amplitude and position of a Gaussian field. A 2D FNO is trained on this data to predict 20-time steps ahead autoregressively from a given initial state. Figure 4 compares the model predictions against ground truth, showing the PRE and confidence bounds from marginal and joint-CP at $90\%$ coverage. The PRE reveals noise artefacts in regions where the field should be zero, highlighting physical inconsistencies despite apparently accurate predictions. Joint-CP bounds are necessarily larger than marginal-CP bounds as they guarantee coverage over the entire spatio-temporal domain rather than individual cells as they span across the spatio-temporal domain as opposed to being cell-wise. As demonstrated in Figure 3, both CP approaches achieve the expected coverage guarantees. Marginal-CP shows linear coverage due to cell-wise averaging, while joint-CP exhibits coverage variations that depend on the calibration dataset (see Appendix C for stability analysis across multiple calibration sets). Additional experimental details are provided in Appendix J.
@@ 5.2 Navier-Stokes Equation
$\overrightarrow{\mathbf{\nabla}}\cdot\overrightarrow{\mathbf{v}}$ $=0$ (11)
@@
$\partialderivative{\overrightarrow{\mathbf{v}}}{t}+(\overrightarrow{\mathbf{v}}\cdot\overrightarrow{\mathbf{\nabla}})\overrightarrow{\mathbf{v}}$ $=\nu\nabla^{2}\overrightarrow{\mathbf{v}}-\nabla P$ (12)
@@
where we are interested in modelling the evolution of the velocity vector $(\overrightarrow{\mathbf{v}}=[u,v])$ and pressure $(P)$ field of an incompressible fluid with kinematic viscosity $(\nu)$. For data generation, Equations 12 and 11 are solved on a domain $x\in[0,1],\;y\in[0,1],\;t\in[0,0.5]$ using a spectral-based solver (Canuto et al., 2007). A 2D multi-variable FNO (Gopakumar et al., 2024b) is trained to model the evolution of velocity and pressure autoregressively up until the $20^{th}$ time instance.
(a) PRE: Momentum Equation
(b) Marginal Bounds
(c) Joint Bounds
Unlike previous examples, the Navier-Stokes case is comprised of two equations. It hence has two PRE estimates: The continuity equation in Equation 11 and the momentum equation Equation 12, representing the conservation of mass and momentum respectively. Our method of performing CP over the residual space using PRE allows us to calibrate the deviation of the model from the physical ground truth concerning each equation. Figure 5 represents the PRE of the momentum equation over the FNO prediction, the upper bounds obtained by performing marginal and joint-CP over the FNO prediction. In Figure 22, the same is depicted for the conservation of mass. Having two PDE residuals provides our framework added scrutiny in identifying relatively inconsistent predictions, as those that violate both bounds can be rejected easily. Further details about the physics, parameterisation of the initial conditions, model and its training can be found in Appendix K. Within the scope of this paper, we limit ourselves to measuring the deviation of the model with the PDE residual. The PRE-CP formulation can be extended to obtain bounds for both the initial and boundary conditions, and this is further explored within Appendix E.
@@ 5.3 Magnetohydrodynamics
$\rho\bigg{(}\partialderivative{\overrightarrow{\mathbf{v}}}{t}+\overrightarrow{\mathbf{v}}\cdot\nabla\overrightarrow{\mathbf{v}}\bigg{)}=\frac{1}{\mu_{0}}\overrightarrow{\mathbf{B}}\times(\overrightarrow{\mathbf{\nabla}}\times\overrightarrow{\mathbf{B}})-\nabla P$ (14)
@@
where the density $(\rho)$, velocity vector $(\overrightarrow{\mathbf{v}}=[u,v])$ and the pressure of plasma is modelled under a magnetic field $(\overrightarrow{\mathbf{B}}=[B_{x},B_{y}])$ across a spatio-temporal domain $x,y\in[0,1]^{2},\;t\in[0,5]$. $\mu_{0}$ is the magnetic permeability of free space. Equations 13, 14, 15, 16 and 17 represent the ideal MHD equations as a combination of the Navier-Stokes equations for fluid flow with Maxwell’s equations of electromagnetism (Alfvén, 1942; Gruber & Rappaz, 1985; Mocz et al., 2014). The equations assume perfect conductivity (no magnetic diffusivity) and no viscosity. We focus our experiment on the modelling of the Orszag-Tang vortex of a turbulent plasma (Orszag & Tang, 1979) with the data being generated using a finite volume method (Eymard et al., 2000). A 2D FNO is trained to model the evolution of all 6 variables over a dataset generated by parameterised initial conditions.
Equations 13, 14, 15, 16 and 17 provide us with five measures of estimating the PRE of the MHD surrogate model. Each PRE estimate depends on a different set of variables associated with the system and allows us to infer errors contributed to each variable accordingly. In Figure 6, slice plots indicating PRE-CP using the induction Equation 16 and the energy Equation 15 is shown for $90\%$ coverage ($\alpha=0.1$). Sliced at $y=0.5m$, the plots indicate the PRE of the mentioned equation along with the marginal and joint bounds across the x-axis for a specific point in time of the simulation. The marginal bounds in both equations provide significantly tighter bounds, whereas the joint bounds provide wider bounds, but their utility comes in identifying the predictions that statistically violate the conservation equations. Plots indicating CP utilising the other residuals (Figures 27 and 28) as well as further details about the physics and the surrogate model can be found in Appendix L.
@@ 5.4 Plasma Modelling within a Tokamak
In (Gopakumar et al., 2024b), the authors model the evolution of plasma blobs within a fusion reactor (known as a tokamak) using an FNO. They explore the case of electrostatic modelling of reduced magnetohydrodynamics with data obtained from the JOREK code (Hoelzl et al., 2021). In the absence of magnetic pressure to confine it, the plasma, driven by kinetic pressure, moves radially outward and collides with the wall of the reactor. The plasma is characterised by density $\rho$, electric potential $\phi$ and Temperature $T$, and the FNO models their spatio-temporal evolution autoregressively. Borrowing upon their pre-trained model and utilising the reduced-MHD equations within the toroidal domain, we demonstrate obtaining calibrated error bars using PRE-CP at scale. The FNO demonstrated in (Gopakumar et al., 2024b) is able to model the plasma six orders of magnitude faster than traditional numerical solvers, and by providing calibrated error bars over the predictions, a wider range of plasma configurations can be validated.
@@
Figure 7: Reduced MHD: PRE-CP using the Temperature equation (Eqn. 3 in (Gopakumar et al., 2024b)) of reduced-MHD to bound the plasma surrogate models. The PRE captures the model error relatively well, allowing us to provide lower and upper error bars corresponding to our required coverage.
We focus on the temperature equation within reduced-MHD (equation 3 within (Gopakumar et al., 2024b)) as it comprises all the variables associated with the plasma. As shown in figure 7, our method can capture the model error across a range of predictions and devise error bars that provide guaranteed coverage without additional data. In figure 7(a), we demonstrate the absolute error in the model prediction of the temperature evolution, correlating that with the PRE over the temperature equations in figure 7(b). By obtaining bounds on the PRE, we can determine the efficacy of the surrogate model in evaluating plasma evolution and determining conditions under which they feel, leading to running JOREK for further evaluation.

### v2

@@ 4.3 Differential Operator: Finite-Difference Stencils as Convolutional Kernels
Calibrating neural PDEs using PRE nonconformity scores requires frequent evaluations of the composite differential operator $D$ in Equation 1. For PDEs, this involves estimating spatio-temporal gradients across the discretised domain, ranging from millions in simple cases to billions of gradient operations for complex physics. To address this computational challenge, we developed a scalable gradient estimation method for evaluating PRE.
We use convolution operations with Finite Difference (FD) stencils as convolutional kernels for gradient estimation (Actor et al., 2020; Chen et al., 2024a, b). For instance, the 2D Laplacian operator $\nabla^{2}$, using a central difference scheme with discretisation $h$, can be approximated by
@@
and used as a kernel. This approach is justified by the mathematical equivalence of FD approximations and discrete convolutions. Both represent matrix-vector multiplications of a block Toeplitz matrix with a field vector (Strang, 1986; Fiorentino & Serra, 1991). The efficiency of this method stems from the optimised implementation of convolution operations in machine learning libraries like PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2015).  [^1] The Basic Linear Algebra Subroutines (BLAS) in these libraries leverage vectorisation and efficient memory access for significant performance gains. Our experiments demonstrate a 1000x speedup using torch.nn.functional.conv3d versus an equivalent numpy finite difference implementation on standard CPU.
The FD approximation offers several advantages over Automatic Differentiation (AD) for our application. It is compatible with CP as a post-hoc measure, requires no architectural modifications, and is model-agnostic. Furthermore, FD implemented via convolutions is more memory-efficient than AD, which requires storing the entire computational graph. As we focus on the (mis)alignment of neural PDEs with Equation 1, we disregard boundary conditions, but demonstrate that they can be accounted for with convolutional padding in Appendix E (Alguacil et al., 2021). Utilising finite difference schemes, the residuals are estimated up to the truncation errors of the Taylor approximation (MacKinnon & Johnson, 1991). Although discretisation plays a role in the width of the error bars, CP-PRE still guarantees coverage, and this is further explored in Section D.1.
@@ 5 Experiments
Figure 3: Validation plots demonstrating coverage guarantee detailed in Equation 5 obtained by performing CP using PRE across experiments. The average empirical coverage obtained experimentally is given on the $y$-axis (ranging from 0 to 1, with 1 representing 100% coverage), while the theoretical coverage is represented on the $x$-axis. We obtain guaranteed coverage while using marginal-CP formulation and near-to-ideal coverage for the joint-CP formulation.
CP-PRE experiments comprise two campaigns. First, we benchmark CP-PRE within standard neural PDEs (Section 5.1 to 5.3). The calibration process (Figure 2) involves: (a) sampling inputs to the model to generate predictions, (b) calculating PRE(s) scores, and (c) calibrating physical error using marginal and joint-CP formulations. Validation uses the same PDE condition bounds as calibration. Within this campaign, we compare our method (CP-PRE) with other methods of providing uncertainty estimation for neural-PDEs. We compare various Bayesian methods along with multivariate inductive conformal prediction to our method as shown in Table 2. We demonstrate that our method is capable of providing valid coverage guarantees for both in and out-of-distribution testing without a significant increase in evaluation times (including time taken for sampling in Bayesian methods, data generation in data-driven CP and subsequent calibration). Methods that attain the required coverage are emboldened. Table 2 provides a qualitative comparison of our method against the other benchmarks and highlights that our method is data-free, requires no modification or sampling and provides guaranteed coverage in a physics-informed manner. We confine our comparison studies in Table 3, 4, and 5 to the Wave, Navier-Stokes and Magnetohydrodynamic equations.
MC Dropout (Gal & Ghahramani, 2016)
Deep Ensemble (Lakshminarayanan et al., 2017)
BNN (MacKay, 1992)
SWA-G (Maddox et al., 2019)
CP-AER (Gopakumar et al., 2024a)
CP-PRE (Ours)
The second campaign (Section 5.4, 5.5) applies CP-PRE to fusion applications. We enhance tokamak plasma behaviour surrogate models to identify erroneous dispersion regions (Section 5.4) and integrate CP-PRE with tokamak design surrogates to identify viable designs and areas needing additional simulations (Section 5.5). This campaign demonstrates the utility of CP-PRE in complex, practical applications. One-dimensional PDE experiments demonstrating CP-PRE are demonstrated in Appendix F and can be reproduced using code in the supplementary material.
@@ 5.1 Wave Equation
Figure 4: Wave: (From left to right) neural PDE (FNO) prediction at the last time instance, physics residual error of the prediction, Upper error bars obtained by performing marginal-CP and joint-CP respectively (90% coverage). For brevity, we have only shown the upper error bars of the symmetric prediction sets. $mod$ represents the modulation function in Table 1. The physical inconsistencies within the residual space of the prediction are calibrated and bounded using CP-PRE.
@@
We solve the field $u$ given in Equation 11 within the domain $x,y\in[-1,1]$, $t\in[0,1.0]$ with wave velocity $c=1.0$ using a spectral solver with periodic boundary conditions (Canuto et al., 2007). The initial conditions are parameterised by the amplitude and position of a Gaussian field. A 2D FNO is trained on this data to predict 20-time steps autoregressively from a given initial state. Figure 4 compares the model predictions against ground truth, showing the PRE and confidence bounds from marginal and joint-CP at $90\%$ coverage. The PRE reveals noise artefacts in regions where the field should be zero, highlighting physical inconsistencies despite apparently accurate predictions. Joint-CP bounds are necessarily larger than marginal-CP bounds as they span across the spatio-temporal domain as opposed to being cell-wise. As demonstrated in Figure 3, both CP approaches achieve the expected coverage guarantees. Table 3 highlights the superior performance of CP-PRE over other methods of UQ, where it guarantees coverage when evaluated for in and out-of-distribution without requiring any additional data, drastically reducing the evaluation time. Marginal-CP shows linear coverage due to cell-wise averaging. At the same time, joint-CP exhibits coverage variations that depend on the calibration dataset (see Table 3 for stability analysis across multiple calibration sets). Additional experimental details are provided in Appendix J.
in-distribution out-of-distribution Time
UQ MSE Coverage $(95\%)$ MSE Coverage $(95\%)$ Train (hr) Eval (s)
Deterministic 1.77e-05 $\pm$ 3.69e-07 - 2.46e-03 $\pm$ 2.00e-05 - 0:38 22
CP-AER 1.76e-05 $\pm$ 4.40e-07 95.70 $\pm$ 0.21 2.46e-03 $\pm$ 1.41e-05 95.59 $\pm$ 0.14 0:38 2022
CP-PRE (Ours) 1.78e-05 $\pm$ 4.61e-07 95.52 $\pm$ 0.21 2.46e-03 $\pm$ 1.25e-05 95.39 $\pm$ 0.12 0:38 32
Table 3: Wave Equation — CP-PRE guarantees coverage across distributions while providing the quickest evaluation time.
@@ 5.2 Navier-Stokes Equation
$\overrightarrow{\mathbf{\nabla}}\cdot\overrightarrow{\mathbf{v}}$ $=0,$ (12)
@@
$\partialderivative{\overrightarrow{\mathbf{v}}}{t}+(\overrightarrow{\mathbf{v}}\cdot\overrightarrow{\mathbf{\nabla}})\overrightarrow{\mathbf{v}}$ $=\nu\nabla^{2}\overrightarrow{\mathbf{v}}-\nabla P,$ (13)
@@
where we are interested in modelling the evolution of the velocity vector $(\overrightarrow{\mathbf{v}}=[u,v])$ and pressure $(P)$ field of an incompressible fluid with kinematic viscosity $(\nu)$. For data generation, Equations 13 and 12 are solved on a domain $x\in[0,1],\;y\in[0,1],\;t\in[0,0.5]$ using a spectral-based solver (Canuto et al., 2007). A 2D multi-variable FNO (Gopakumar et al., 2024b) is trained to model the evolution of velocity and pressure autoregressively up until the $20^{th}$ time instance.
in-distribution out-of-distribution Time
UQ MSE Coverage $(95\%)$ MSE Coverage $(95\%)$ Train (hr) Eval (s)
Deterministic 1.05e-04 $\pm$ 6.91e-06 - 3.67e-03 $\pm$ 5.30e-05 - 3:22 25
CP-AER 1.05e-04 $\pm$ 6.58e-06 95.56 $\pm$ 0.40 3.66e-03 $\pm$ 2.81e-05 95.54 $\pm$ 0.15 3:22 20026
CP-PRE (Ours) 1.07e-04 $\pm$ 5.18e-06 95.44 $\pm$ 0.22 3.70e-03 $\pm$ 4.23e-05 95.57 $\pm$ 0.14 3:22 134
Table 4: Navier-Stokes Equations — CP-PRE guarantees coverage across distributions while providing the quickest evaluation time.
Unlike the previous example, the Navier-Stokes case has two equations with corresponding PRE estimates: continuity (Equation 12) and momentum (Equation 13) for mass and momentum conservation. Our CP-PRE method calibrates model deviation from physical ground truth for each equation. Table 4 shows that Bayesian methods fail to provide coverage for FNO modelling, while multivariate CP (CP-AER) provides coverage but requires substantially higher evaluation time arising from the generation of simulation data for calibration. Figures 20 and 21 show PRE bounds from both equations. Having two PDE residuals enables easier rejection of predictions violating both bounds. Physics details, parameterisation, and training are in Appendix K. Within the scope of this paper, we limit ourselves to measuring the deviation of the model with the PDE residual. The CP-PRE formulation can be extended to obtain bounds for both the initial and boundary conditions, further explored within Appendix E.
@@ 5.3 Magnetohydrodynamics
$\rho\bigg{(}\partialderivative{\overrightarrow{\mathbf{v}}}{t}+\overrightarrow{\mathbf{v}}\cdot\nabla\overrightarrow{\mathbf{v}}\bigg{)}=\frac{1}{\mu_{0}}\overrightarrow{\mathbf{B}}\times(\overrightarrow{\mathbf{\nabla}}\times\overrightarrow{\mathbf{B}})-\nabla P,$ (15)
@@
where the density $(\rho)$, velocity vector $(\overrightarrow{\mathbf{v}}=[u,v])$ and the pressure of plasma is modelled under a magnetic field $(\overrightarrow{\mathbf{B}}=[B_{x},B_{y}])$ across a spatio-temporal domain $x,y\in[0,1]^{2},\;t\in[0,5]$. $\mu_{0}$ is the magnetic permeability of free space. Equations 14, 15, 16, 17 and 18 represent the ideal MHD equations as a combination of the Navier-Stokes equations for fluid flow with Maxwell’s equations of electromagnetism (Alfvén, 1942; Gruber & Rappaz, 1985; Mocz et al., 2014). The equations assume perfect conductivity (no magnetic diffusivity) and no viscosity. We focus our experiment on the modelling of the Orszag-Tang vortex of a turbulent plasma (Orszag & Tang, 1979) with the data being generated using a finite volume method (Eymard et al., 2000). A 2D FNO is trained to model the evolution of all 6 variables over a dataset generated by parameterised initial conditions.
in-distribution out-of-distribution Time
UQ MSE Coverage $(95\%)$ MSE Coverage $(95\%)$ Train (hr) Eval (s)
Deterministic 2.20e-03 $\pm$ 5.20e-03 - 4.71e-02 $\pm$ 1.06e-03 - 5:00 40
CP-AER 2.20e-03 $\pm$ 4.38e-05 95.61 $\pm$ 0.26 4.69e-02 $\pm$ 8.18e-04 95.60 $\pm$ 0.27 5:00 40042
CP-PRE (Ours) 2.20e-03 $\pm$ 4.96e-03 95.54 $\pm$ 0.18 4.71e-02 $\pm$ 1.06e-03 95.67 $\pm$ 0.22 5:00 482
Table 5: Magnetohydrodynamic Equations — CP-PRE guarantees coverage across distributions with only a marginal increase in evaluation time, arising from residual evaluation over a larger family of PDEs (see Equations 14, 15, 16, 17 and 18).
Equations 14, 15, 16, 17 and 18 provide us with five measures of estimating the PRE of the MHD surrogate model. Each PRE estimate depends on a different set of variables associated with the system, allowing us to infer errors contributed to each variable accordingly. Table 5 demonstrates that the Bayesian methods fail miserably in providing valid error bars over the MHD case. The CP-AER provides guaranteed coverage but is heavily dependent on simulation data, adding to the computational expense. CP-PRE using induction (Equation 17) and energy (Equation 16) are shown for $90\%$ coverage $(\alpha=0.1)$, sliced at $y=0.5m$. The plots show PRE along the x-axis with marginal and joint bounds at a specific simulation time. Marginal bounds are significantly tighter, while joint bounds are wider but useful for identifying predictions that violate conservation equations. Additional CP plots using other residuals (Figures 27 and 28) and physics/surrogate model details are in Appendix L.
@@ 5.4 Plasma Modelling within a Tokamak
In (Gopakumar et al., 2024b), the authors model the evolution of plasma blobs within a fusion reactor (known as a tokamak) using an FNO. They explore the case of electrostatic modelling of reduced magnetohydrodynamics with data obtained from the JOREK code (Hoelzl et al., 2021). In the absence of magnetic pressure to confine it, the plasma, driven by kinetic pressure, moves radially outward and collides with the wall of the reactor. The plasma is characterised by density $\rho$, electric potential $\phi$ and Temperature $T$, and the FNO models their spatio-temporal evolution autoregressively. Borrowing that pre-trained model and utilising the reduced-MHD equations within the toroidal domain, we demonstrate obtaining calibrated error bars using CP-PRE at scale. The FNO demonstrated in (Gopakumar et al., 2024b) can model the plasma six orders of magnitude faster than traditional numerical solvers, and by providing calibrated error bars over the predictions, a wider range of plasma configurations can be validated.
@@
Figure 5: Reduced MHD: CP-PRE using the Temperature equation (Eqn. 3 in (Gopakumar et al., 2024b)) of reduced-MHD to bound the plasma surrogate models. The PRE captures the model error relatively well, allowing us to provide lower and upper error bars corresponding to our required coverage.
We focus on the temperature equation within reduced-MHD (equation 3 within (Gopakumar et al., 2024b)) as it comprises all the variables associated with the plasma. As shown in Figure 5, our method can capture the model error across a range of predictions and devise error bars that provide guaranteed coverage without additional data. In figure 5(a), we demonstrate the absolute error in the model prediction of the temperature evolution, correlating that with the PRE over the temperature equations in figure 5(b). By obtaining bounds on the PRE, we can determine the efficacy of the surrogate model in evaluating plasma evolution and identify conditions under which it fails, running the MHD code JOREK under those failed conditions.